# The Dynamical History of the Kepler-221 Planet System

### v1

@@ Appendix C Mass Dependence in Phase IV
$a_{\mathrm{cde}}=\frac{\Delta({P_{\mathrm{e}}/P_{\mathrm{d}}})}{\Delta(P_{\mathrm{d}}/P_{\mathrm{c}})}=-\frac{P_{\mathrm{c}}^{2}}{P_{\mathrm{d}}^{2}}\frac{\Delta P_{\mathrm{e}}}{\Delta P_{\mathrm{c}}}$ (20)
@@
$\frac{P_{\mathrm{3}}}{P_{\mathrm{2}}}=\frac{qP_{1}}{rP_{1}-pP_{2}}$ (21)
The total angular momentum for the three planets in 3BR is conserved and can be expressed as:
$L_{\mathrm{tot}}=\left(\frac{G^{2}M^{2}}{2\pi}\right)^{1/3}\left(m_{\mathrm{1}}P_{\mathrm{1}}^{1/3}+m_{\mathrm{2}}P_{\mathrm{2}}^{1/3}+m_{\mathrm{3}}P_{\mathrm{3}}^{1/3}\right)$ (22)
@@
$L_{\mathrm{tot}}=\left(\frac{G^{2}M^{2}}{2\pi}\right)^{1/3}\left[m_{1}\left(\frac{pP_{2}P_{3}}{rP_{3}-qP_{2}}\right)^{1/3}+m_{2}P_{2}^{1/3}+m_{3}P_{3}^{1/3}\right]$ (23)
@@
$\frac{m_{2}}{P_{2}^{2/3}}+\frac{m_{3}}{P_{3}^{2/3}}Y_{32}+\frac{p(rP_{3}^{2}-qP_{2}^{2}Y_{32})m_{1}}{(qP_{2}-rP_{3})^{2}\left(\frac{pP_{2}P_{3}}{rP_{3}-qP_{2}}\right)^{2/3}}=0$ (24)
@@
$Y_{32}=\frac{p^{1/3}rP_{\mathrm{32}}^{2}m_{1}+\left(rP_{\mathrm{32}}-q\right)^{4/3}P_{\mathrm{32}}^{2/3}m_{2}}{p^{1/3}qm_{1}-\left(rP_{\mathrm{32}}-q\right)^{4/3}m_{3}}$ (25)
@@
$Y_{21}=\frac{q^{1/3}pP_{\mathrm{21}}^{2}m_{3}-\left(r-pP_{\mathrm{21}}\right)^{4/3}P_{\mathrm{21}}^{2/3}m_{2}}{q^{1/3}rm_{3}+\left(r-pP_{\mathrm{21}}\right)^{4/3}m_{2}}$ (26)
@@
$\frac{\Delta P_{\mathrm{e}}}{\Delta P_{\mathrm{c}}}=\frac{15m_{\mathrm{b}}+15.12m_{\mathrm{c}}}{m_{\mathrm{b}}-7.27m_{\mathrm{e}}}.$ (27)
@@
$a_{\mathrm{cde}}=\frac{4.90m_{\mathrm{b}}+4.94m_{\mathrm{c}}}{7.27m_{\mathrm{e}}-m_{\mathrm{b}}}$ (28)
@@ Appendix D Origin of the Migration of Planet d in Phase IV
For the blue line in Figure 14, planet d migrates neither outward nor inward because no tidal damping is applied on planet d. For the yellow line, tidal damping is added on planets b, c, and e but not on planet d, and planet d would migrate outward at the speed faster for more massive planet d. This is because the tidal damping on planets b, c, and e would expand the 3BR, increasing the relative period ratio. During this process, tidal dissipation conserves angular momentum, planets b and c would migrate inward and planets d and e would migrate outward slightly due to the angular momentum exchange. A more massive d planet will increase the eccentricity of the other planets, increasing their relative tidal dissipation and the outward migration speed of planet d. Such effect of outward migration explains the difference between the black line and the colored line in panel b of Figure 5, in which the simulation (colored line) deviates from the analytical result (black line) because planet d also moves outward during the orbital expansion phase.

### v2

@@ Appendix C Mass dependence in Phase IV
$a_{\mathrm{cde}}=\frac{\Delta({P_{\mathrm{e}}/P_{\mathrm{d}}})}{\Delta(P_{\mathrm{d}}/P_{\mathrm{c}})}=-\frac{P_{\mathrm{c}}^{2}}{P_{\mathrm{d}}^{2}}\frac{\Delta P_{\mathrm{e}}}{\Delta P_{\mathrm{c}}},$ (20)
@@
$\frac{P_{\mathrm{3}}}{P_{\mathrm{2}}}=\frac{qP_{1}}{rP_{1}-pP_{2}},$ (21)
The total angular momentum for the three planets in a 3BR is conserved and can be expressed as:
$L_{\mathrm{tot}}=\left(\frac{G^{2}M^{2}}{2\pi}\right)^{1/3}\left(m_{\mathrm{1}}P_{\mathrm{1}}^{1/3}+m_{\mathrm{2}}P_{\mathrm{2}}^{1/3}+m_{\mathrm{3}}P_{\mathrm{3}}^{1/3}\right),$ (22)
@@
$L_{\mathrm{tot}}=\left(\frac{G^{2}M^{2}}{2\pi}\right)^{1/3}\left[m_{1}\left(\frac{pP_{2}P_{3}}{rP_{3}-qP_{2}}\right)^{1/3}+m_{2}P_{2}^{1/3}+m_{3}P_{3}^{1/3}\right],$ (23)
@@
$\frac{m_{2}}{P_{2}^{2/3}}+\frac{m_{3}}{P_{3}^{2/3}}Y_{32}+\frac{p(rP_{3}^{2}-qP_{2}^{2}Y_{32})m_{1}}{(qP_{2}-rP_{3})^{2}\left(\frac{pP_{2}P_{3}}{rP_{3}-qP_{2}}\right)^{2/3}}=0,$ (24)
@@
$Y_{32}=\frac{p^{1/3}rP_{\mathrm{32}}^{2}m_{1}+\left(rP_{\mathrm{32}}-q\right)^{4/3}P_{\mathrm{32}}^{2/3}m_{2}}{p^{1/3}qm_{1}-\left(rP_{\mathrm{32}}-q\right)^{4/3}m_{3}},$ (25)
@@
$Y_{21}=\frac{q^{1/3}pP_{\mathrm{21}}^{2}m_{3}-\left(r-pP_{\mathrm{21}}\right)^{4/3}P_{\mathrm{21}}^{2/3}m_{2}}{q^{1/3}rm_{3}+\left(r-pP_{\mathrm{21}}\right)^{4/3}m_{2}},$ (26)
@@
$\frac{\Delta P_{\mathrm{e}}}{\Delta P_{\mathrm{c}}}=\frac{15m_{\mathrm{b}}+15.12m_{\mathrm{c}}}{m_{\mathrm{b}}-7.27m_{\mathrm{e}}},$ (27)
@@
$a_{\mathrm{cde}}=\frac{4.90m_{\mathrm{b}}+4.94m_{\mathrm{c}}}{7.27m_{\mathrm{e}}-m_{\mathrm{b}}},$ (28)
@@ Appendix D Origin of the migration of planet d in Phase IV
For the blue line in Figure 14, planet d migrates neither outward nor inward because no tidal damping is applied on planet d. For the yellow line, tidal damping is added on planets b, c, and e but not on planet d, and planet d would migrate outward at a faster speed with a more massive planet d. This is because the tidal damping on planets b, c, and e would expand the 3BR, increasing the relative period ratio. During this process, tidal dissipation conserves angular momentum, planets b and c would migrate inward and planets d and e would migrate outward slightly due to the angular momentum exchange. A more massive d planet will increase the eccentricity of the other planets, increasing their relative tidal dissipation and the outward migration speed of planet d. Such effect of outward migration explains the difference between the black line and the colored line in panel b of Figure 5, in which the simulation (colored line) deviates from the analytical result (black line) because planet d also moves outward during the orbital expansion phase.